OpenStudy (anonymous):

Which of the point-slope equations below is correct for the line that passes through points (-1,-4) and (3,2)? Check all that apply.

OpenStudy (anonymous):

first find the slope of the line

OpenStudy (anonymous):

which is (y2-y1)/(x2-x1)

OpenStudy (anonymous):

I dont even know how how to find the aslope...

OpenStudy (anonymous):

then choose any point, plug in the x and y values into the point slope format and ur done

OpenStudy (anonymous):

(-1,-4) and (3, 2)

OpenStudy (anonymous):

y + 4 = 2/3(x + 1) B. y -1 = 3/2(x + 3) C. y - 2 = 3/2(x - 3) D. y + 4 = 3/2(x - 1) E. y + 4 = 3/2(x + 1)

OpenStudy (anonymous):

(y2-y1) and (x2-x1)

OpenStudy (anonymous):

those are the answer choices...

OpenStudy (anonymous):

u wanna find the slope? lol

OpenStudy (anonymous):

I DONT KNOW HOW TO EVEN FIND THE SLOPE